OpenStudy (anonymous):

betrice decided to sell some of her dolls in her doll colection. she sold 20 dolls. this way 40% of her collection. how many dolls did she have left?

OpenStudy (anonymous):

40%/100% For every 20 dolls its 40% 40+40= 80%, and you need 20% more to get 100%. Half of 40% is 20%, and half of 20 dolls is 10. 10 dolls =20% 80% + 20%= 100% 40 + 10 = 50 50 dolls=100% But since she sold 20 dolls already... 50-20= 30 This means she had 30 dolls left to sell.